OpenStudy (anonymous):

What is the solution of the system of inequalities? y≥x2 +2x+2 y<−x2 −4x−2

OpenStudy (helder_edwin):

u have \[ \large y\geq x^2+2x+2 \] \[ \large y<-x^2-4x-2 \] complete the squares to be able to sketch the parabolas.

OpenStudy (anonymous):

How do you complete the square?

OpenStudy (helder_edwin):

\[ \large y-2+1\geq x^2+2x+1 \] \[ \large y-1\geq(x+1)^2 \] forget for a moment u r dealing with an inequality and graph \[ \large y-1=(x+1)^2 \] do the same with the other one.